two men A and B are standing on opposite edge of a 6m long platform which is further kept on smooth floor.they start moving towards each other and finally meet at the midpoint of platform.find the displacement of platform if masses of A B and platform are 40kg,60kg,50 kg respectively

Grade:11

1 Answers

Vikas TU
7 years ago
m1x1 = m2x2
therfore,
60x +  50(6 -x ) =  40(6 +x)
60x + 300 – 50x = 240 + 40x
30x = 60
x = 2 meters.
 
Displacemnt of Platform => 6-x => 6-2 = 4 meters